two angles are supplementary. the measure of the larger angle is four times the measure of the smaller angle. find the measure of each angle.
Answer by oberobic(2304)

Two angles are supplementary.
x + y = 180 degrees
.
x = 4y
.
Substitute
.
4y + y = 180
5y = 180
y = 36
.
4y = 4*36 = 144
.
Check
36+144 = 180
.
Answer: The two angles are 36 and 144 degrees.